OpenStudy (anonymous):

what is the slope of the line that passes through the points (1 17) and (4 2)

OpenStudy (scarlettfarra2000):

(17+2)(1+4)=15/3=5

OpenStudy (scarlettfarra2000):

Also you probably should have put this in Mathematics someone would have helped you sooner

OpenStudy (anonymous):

oh srry didnt see that